# Exploring the Boundaries of On-Device Inference: When Tiny Falls Short, Go Hierarchical

### v1

@@
Abstract.
On-device inference holds great potential for increased energy efficiency, responsiveness, and privacy in edge ML systems. However, due to less capable ML models that can be embedded in resource-limited devices, use cases are limited to simple inference tasks such as visual keyword spotting, gesture recognition, and predictive analytics. In this context, the Hierarchical Inference (HI) system has emerged as a promising solution that augments the capabilities of the local ML by offloading selected samples to an edge server/cloud for remote ML inference. Existing works demonstrate through simulation that HI improves accuracy. However, they do not account for the latency and energy consumption on the device, nor do they consider three key heterogeneous dimensions that characterize ML systems: hardware, network connectivity, and models.
In contrast, this paper systematically compares the performance of HI with on-device inference based on measurements of accuracy, latency, and energy for running embedded ML models on five devices with different capabilities and three image classification datasets. For a given accuracy requirement, the HI systems we designed achieved up to $73$% lower latency and up to $77$% lower device energy consumption than an on-device inference system. The key to building an efficient HI system is the availability of small-size, reasonably accurate on-device models whose outputs can be effectively differentiated for samples that require remote inference. Despite the performance gains, HI requires on-device inference for all samples, which adds a fixed overhead to its latency and energy consumption. Therefore, we design a hybrid system, Early Exit with HI (EE-HI), and demonstrate that compared to HI, EE-HI reduces the latency up to $59.7$% and lowers the device’s energy consumption up to $60.4$%.
1. Introduction
@@
Deep Learning (DL) inference on data generated by end devices such as IoT sensors, smartphones, and drones enables significant improvements in operational efficiency and functionality across various sectors, including industrial automation, smart cities, remote healthcare, smart agriculture, and other sectors. Performing DL inference locally on the device has the added advantage of increased energy efficiency (Banbury et al., 2021b), responsiveness, and privacy and holds great promise for real-time decision-making. Thus, significant efforts have been directed towards designing and developing compact DL models specifically tailored for deployment on end devices (Deng et al., 2020; Chen et al., 2020; Sanchez-Iborra and Skarmeta, 2020). This resulted in a plethora of DL models, such as MobileNet (Sandler et al., 2018), EfficientNet (Tan and Le, 2019), and Gemma 2B large language model (DeepMind), 2024), which can be embedded on moderately powerful smartphones. Additionally, tinyML models like ResNet-8 (Banbury et al., 2021a) have been designed for extremely resource-limited micro-controller units (MCUs).
@@
Figure 1. Measured inference accuracy and latency per inference for various models on different MCUs: Arduino Nano, ESP-32, and Coral Micro, for the MNIST, CIFAR-10, and ImageNet-1K datasets. As the inference complexity increases from MNIST to ImageNet-1K, even resorting to larger DL models and powerful MCU Coral Micro (with TPU) may be insufficient to satisfy QoS requirements; for example $90$% accuracy (green line) and $250$ ms response time (red line).
@@
Especially when considering DL inference on MCUs, despite the advances in models’ optimizations and improvements, the end-product use cases of on-device inference are limited to applications involving simple inference tasks – visual keyword spotting on smartphones, gesture recognition on smart cameras, and predictive analytics for industrial machines on sensors (imagimob, 2024). We indeed argue that on-device inference becomes inadequate as the complexity of the inference tasks increases. For this purpose, consider an image classification application [^1] We focus on image classification datasets and models as they span a broad range of complexity and are widely used benchmarks for evaluating the performance of DL systems (Reddi et al., 2020). with a QoS requirement of at least $90$% accuracy and maximum $250$ ms latency [^2] This value is motivated by the latency requirement of $100$ to $500$ ms in real-time computer vision applications such as scene text detection (Zhou et al., 2017), and image captioning (Zhou et al., 2021).. In Fig. 1, we present the accuracy and latency for different ML models on three MCUs: Arduino Nano, ESP-32, and Coral Micro (cf. Table 1). Of the MCUs, Coral Micro is the most powerful, featuring a Tensor Processing Unit (TPU). For the inference tasks, we categorize MNIST image classification as a simple task, CIFAR-10 as a moderately complex task, and ImageNet-1K as a complex task. Observe that Arduino Nano and ESP-32 satisfy the QoS requirement for the MNIST dataset using Logistic Regression (LR). But both devices fall short of latency and accuracy requirements for CIFAR-10 dataset because running the state-of-the-art tinyML model ResNet-8, which has $87$% accuracy, takes at least $600$ ms. Coral Micro can meet the QoS requirements for CIFAR-10, using a bigger DL model ResNet-56, but it cannot do so for ImageNet-1K.
The issues discussed above hinder the large-scale adoption of on-device inference. An alternate solution for devices connected to a network is to use an edge server/cloud that uses a state-of-the-art large-size DL model. By offloading all the samples for remote inference on the large-size DL model, a device can achieve higher accuracy rather than inferring locally. However, fully offloading ML inference execution to more powerful computing devices risks not leveraging on-device capabilities due to additional network latency and connectivity dependency, forgoing the benefits of increased energy efficiency and responsiveness of on-device inference.
@@
In this context, Hierarchical Inference (HI) (Moothedath et al., 2024) has recently emerged as a novel distributed ML inference technique that mitigates the limitations of on-device inference by using remote inference when necessary. Under a HI system, inference is performed first on the local DL model for each data sample. A data sample for which the local inference is likely correct is a simple data sample, and its inference is accepted, while a data sample for which the local inference is likely incorrect is a complex data sample, which is offloaded. By only offloading the complex samples, HI reaps the benefits of performing inference on end devices without compromising on accuracy, and thus, it received considerable attention in the recent past (Nikoloska and Zlatanov, 2021; Al-Atat et al., 2023; Behera et al., 2023; Moothedath et al., 2024; Beytur et al., 2024).
@@
HI operates through a decision algorithm that acts on the on-device (local ML) model output for each input sample and decides whether to accept the local inference or offload the sample. Consequently, HI requires on-device inference for every sample, introducing a fixed overhead to latency and energy consumption per inference. However, all the existing work on HI only studied the accuracy improvement over the on-device inference without accounting for its latency and on-device energy consumption overheads. Further, their conclusions were based only on simulations using abstract costs for offloading and misclassification errors. Additionally, they ignore the diversity and heterogeneity of edge ML systems, including varying hardware, network connectivity, and models.
@@
In this paper, we address the above shortcomings and make multiple contributions. We perform accuracy, latency, and energy measurements for running embedded DL models on five devices with different capabilities and three image classification datasets (please refer to Section 3 for additional details about the devices and models used). These measurements will be of independent interest to the embedded ML community. Using the measurements, we empirically evaluate the performance of HI and on-device inference and demonstrate that HI improves accuracy but has higher latency and energy consumption. This raises the question of the true benefits of using HI. We answer this question in light of QoS requirements. We demonstrate that a HI system achieves a given accuracy requirement using a small on-device model, even if this model’s accuracy is below the specified requirement. Consequently, the HI system using the small model achieves lower latency and energy consumption than a larger on-device model that meets the initial accuracy requirement. This establishes the benefit of using HI over on-device inference. For instance, on the Raspberry Pi, with an accuracy requirement of $90$% for CIFAR-10 image classification, we demonstrate that using a smaller model, ResNet-8, with HI achieves $95.9$% accuracy and has $73$% lower latency and $77$% lower energy consumption than the on-device larger model, ResNet-56, with $93.7$% accuracy. Additionally, HI achieves up to $11$% (absolute difference) higher accuracy for a given latency and energy requirement for the systems we studied.
However, for ImageNet-1K image classification, offloading for remote inference remains the only viable solution on Coral Micro and Raspberry Pi due to the lack of on-device models that can provide the necessary lower latency and energy consumption. To alleviate this issue, we resort to the early exit technique (Teerapittayanon et al., 2016), which improves HI by reducing the latency of the on-device inference. Using this technique, we design a hybrid system, Early Exit with HI (EE-HI), and demonstrate that compared to HI, EE-HI reduces the latency up to $59.7$% and the device’s energy consumption up to $60.4$%.
@@
The rest of the paper is organized as follows. In Section 2, we discuss the existing edge ML inference strategies and the related work. In Section 3, we describe the system methodology used for measurements. We present the measurement results and compare the performance of HI, on-device, and remote inference in Section 4. We present the design and the results of EE-HI in Section 5 and conclude in Section 6.
2. Existing Strategies for Edge ML Inference
@@
In this section, we discuss the related works in edge ML inference systems. The methods that use on-device inference can be broadly classified into 1) inference load balancing, 2) DNN partitioning, 3) on-device inference, and 4) HI. In the following, an inference task refers to outputting the inference for a given data sample at the device. For the on-device inference strategy, this task is to run the local DL model on the data sample. For remote ML, this task involves offloading the data sample for remote ML inference and receiving the results. The HI this task involves using standalone on-device inference or on-device in conjunction with remote inference based on the HI offloading decision.
Inference Load Balancing: Since the initial proposal of edge computing (Satyanarayanan et al., 2009), significant attention has been devoted to the computational offloading for generic compute-intensive applications (Mach and Becvar, 2017). Due to the growing importance of edge intelligence, recent studies have examined computation offloading for applications using ML inference (Wang, 2018; Ogden and Guo, 2020; Fresa and Champati, 2023). This line of research aims to load-balance the inference task by partitioning the data samples between the device and the server, considering parameters such as job execution times, communication times, energy consumption, and average test accuracy of ML models. However, in (Al-Atat et al., 2023), Ghina et al. demonstrated that load-balancing data samples may lead to adversarial scenarios where most data samples scheduled on the device receive incorrect inferences. HI circumvents this issue by utilizing the local DL output and offloading a data sample if its local inference is likely incorrect.
DNN Partitioning: DNN partitioning, proposed in (Kang et al., 2017), partitions the layers of a large-size DL model and deploys the front layers on the mobile device while the deep layers are deployed on an edge server. It received considerable attention recently (Li et al., 2020; Hu and Li, 2022). However, as demonstrated in (Al-Atat et al., 2023), for DNN partitioning to be beneficial in reducing the execution time per inference, the processing times of the DL layers on the mobile device should be small relative to the communication time of the data generated between layers. Thus, this technique requires mobile GPUs, making it infeasible for resource-constrained end devices such as IoT devices.
@@
On-Device Inference: As discussed in Section 1, considerable effort has been dedicated to the design of compact DL models tailored for deployment on end devices (Deng et al., 2020; Chen et al., 2020; Sanchez-Iborra and Skarmeta, 2020). Further, techniques such as early exit (Teerapittayanon et al., 2016) have been explored to reduce the execution time of DL inference. As a result, state-of-the-art compact DL models can perform simple ML inference tasks on devices, ranging from face recognition on mobile phones (Lane et al., 2017) to visual wake word and keyword spotting on IoT devices (Banbury et al., 2021a). However, the main focus in designing these models has been improving accuracy, while time per inference and energy consumption have not been critically studied.
To understand the different efficiency aspects of ML models, their performance is required to be measured. However, the complexity of measuring ML performance on resource-constrained devices is challenging. To alleviate such a challenge, an ML performance (MLPerf) benchmarking organization has recently proposed a benchmarking suite, TinyMLPerf, for TinyML models to perform simple inference tasks (Banbury et al., 2021a). Throughput and latency measurements were presented in (Torelli and Bangale, 2021) for ResNet-50 and MobileNet on Jetson Nano and Coral Micro. More fine-grained and precise profiling on Android devices has been proposed by Chu et al. (Chu et al., 2023), who present nnPerf, a profiler for TFLite models. However, none of the above works provide energy measurements or compare edge ML strategies.
@@
Hierarchical Inference: The key to implementing HI is the HI decision module/algorithm for differentiating simple and complex samples and making the offloading decision. To this end, all existing works use the soft-max value (confidence) output by the local DL model corresponding to each class. An image is classified into the class that has the maximum soft-max value. Nikoloska and Zlatanov (Nikoloska and Zlatanov, 2021) computed the HI offloading decision by computing a threshold for the maximum soft-max value based on the transmission energy constraint of the device. Al-Atat et al. (Al-Atat et al., 2023) proposed a general definition for HI, presented multiple use cases, and compared HI with existing distributed DL inference approaches at the edge. Similar to (Nikoloska and Zlatanov, 2021), a threshold was computed based on the trade-off between local misclassifications and offloading costs. Behera et al. (Behera et al., 2023) showed that using a binary LR on the first two highest soft-max values can improve the HI offloading decision. The binary LR is trained to learn which images are correctly classified and which are incorrectly classified by the local DL. This paper uses the binary LR as the HI decision algorithm. All the above works, however, use simulation to evaluate their performance against on-device inference using abstract costs for offloading and misclassification errors.
2.1. Early Exit with Hierarchical Inference
As stated in Section 1, lower-latency on-device models are necessary for HI implementation. Toward this end, we propose the EE-HI system by combining the HI system with the proven inference acceleration technique early exit (Teerapittayanon et al., 2016). To this end, we modify the base models by adding early exit branches to create new local ML models. This allows for inferences to be accepted early based on the likelihood of the data sample being a simple sample. After the data samples are processed by the early exit local ML, HI is applied on top of it. A binary LR model decides if the local inference (at the final layer) for each data sample suffices or if it needs to be offloaded to the remote server for further inference. This method provides a dynamic and adaptive approach to handling a variety of samples efficiently, optimizing both time and resource utilization. More details of this method are presented in Section 5.
3. System and Methodology
3.1. Devices

### v2

@@
Abstract
On-device inference offers significant benefits in edge ML systems, such as improved energy efficiency, responsiveness, and privacy, compared to traditional centralized approaches. However, the resource constraints of embedded devices limit their use to simple inference tasks, creating a trade-off between efficiency and capability. In this context, the Hierarchical Inference (HI) system has emerged as a promising solution that augments the capabilities of the local ML by offloading selected samples to an edge server/cloud for remote ML inference. Existing works, primarily based on simulations, demonstrate that HI improves accuracy. However, they fail to account for the latency and energy consumption in real-world deployments, nor do they consider three key heterogeneous components that characterize ML-enabled IoT systems: hardware, network connectivity, and models. To bridge this gap, this paper systematically evaluates HI against standalone on-device inference by analyzing accuracy, latency, and energy trade-offs across five devices and three image classification datasets. Our findings show that, for a given accuracy requirement, the HI approach we designed achieved up to $73$% lower latency and up to $77$% lower device energy consumption than an on-device inference system. Despite these gains, HI introduces a fixed energy and latency overhead from on-device inference for all samples. To address this, we propose a hybrid system called Early Exit with HI (EE-HI) and demonstrate that, compared to HI, EE-HI reduces the latency up to $59.7$% and lowers the device’s energy consumption up to $60.4$%. These findings demonstrate the potential of HI and EE-HI to enable more efficient ML in IoT systems.
Index Terms:
Machine learning, on-device inference, TinyML, Hierarchical Inference, Early Exit, processing time and energy measurements.
I Introduction
@@
Deep Learning (DL) inference on data generated by end devices such as IoT sensors, smartphones, and drones enables significant improvements in operational efficiency and functionality across various sectors, including industrial automation, smart cities, remote healthcare, smart agriculture, and other sectors. Compared to offloading to the cloud, performing DL inference locally on an IoT device has the added advantage of increased energy efficiency [1], responsiveness, and privacy and holds great promise for real-time decision-making. Thus, significant efforts have been directed towards designing and developing compact DL models tailored explicitly for deployment on end devices [2, 3, 4]. This resulted in a plethora of DL models, such as MobileNet [5], EfficientNet [6], and Gemma 2B large language model [7], which can be embedded on moderately powerful smartphones. TinyML models like ResNet-8 [8] have also been designed for extremely resource-limited IoT sensors with micro-controller units (MCUs).
@@
Figure 1: Measured inference accuracy and latency per inference for various models on different MCUs: Arduino Nano, ESP-32, and Coral Micro, for the MNIST, CIFAR-10, and ImageNet-1K datasets. As the inference complexity increases from MNIST to ImageNet-1K, even resorting to larger DL models and powerful MCU Coral Micro (with TPU) may be insufficient to satisfy QoS requirements: $90$% accuracy (green line) and $250$ ms response time (red line).
@@
Despite the advances in ML model optimization, the large-scale adoption of on-device inference is limited to applications involving simple inference tasks such as visual keyword spotting on smartphones, gesture recognition on smart cameras, and predictive analytics for industrial machines on IoT sensors [9]. We indeed argue that on-device inference becomes inadequate as the complexity of the inference tasks increases. For this purpose, consider an image classification application [^1] We focus on image classification datasets and models as they span a broad range of complexity and are widely used benchmarks for evaluating the performance of DL systems [10]. with a QoS requirement of at least $90$% accuracy and maximum $250$ ms latency [^2] This value is motivated by the latency requirement of $100$ to $500$ ms in real-time computer vision applications such as scene text detection [11], and image captioning [12].. In Fig. 1, we present the accuracy and latency for different ML models on three MCUs: Arduino Nano, ESP-32, and Coral Micro (cf. Table I). Of the MCUs, Coral Micro is the most powerful, featuring a Tensor Processing Unit (TPU). For the inference tasks, we categorize MNIST image classification as a simple task, CIFAR-10 as a moderately complex task, and ImageNet-1K as a complex task. Observe that Arduino Nano and ESP-32 satisfy the QoS requirement for the MNIST dataset using Logistic Regression (LR). But both devices fall short of latency and accuracy requirements for CIFAR-10 dataset because running the state-of-the-art tinyML model ResNet-8, which has $87$% accuracy, takes at least $600$ ms. Coral Micro can meet the QoS requirements for CIFAR-10, using a bigger DL model ResNet-56, but it cannot do so for ImageNet-1K. Thus, resorting to more computationally capable end devices, in the microcontroller space, may still be inadequate to support emerging ML applications with ever-increasing computation demands.
An alternative solution for network-connected devices is to leverage an edge server or cloud equipped with a state-of-the-art, large-scale deep learning model. By offloading all the data samples for remote inference on the large-size DL model, a device can achieve higher accuracy rather than inferring locally. However, fully offloading the execution of the DL inference risks: (i) underutilizing the on-device capabilities, (ii) introducing additional network latency and connectivity dependency, and (iii) forfeiting the energy efficiency and responsiveness benefits of the on-device inference. Consequently, extensive research has lately focused on distributed inference techniques that use both on-device and remote-server computing capacity. These techniques include (i) DNN-partitioning [13], (ii) inference load balancing [14], and (iii) Hierarchical Inference (HI) [15]. DNN-partitioning computes inference using a large-size DL model by partitioning the model’s layers between the device and a remote server. Nevertheless, the benefits of this technique have only been realized for computationally powerful devices (such as smartphones) with mobile GPUs [16, 17]. Furthermore, the technique is not feasible for proprietary DL models, as the model architecture may not be openly available. The inference load-balancing approach divides the computational load between the device and the server to minimize inference time or energy consumption [18, 19, 14].
Figure 2: HI framework for DL inference at the edge.
@@
HI [15] paradigm shown in Fig. 2, mitigates the limitations of on-device inference by using remote inference only when on-device inference is likely incorrect. A data sample for which the on-device inference is likely correct is a simple data sample, and its inference is accepted. In contrast, a data sample deemed likely to result in incorrect local inference is considered a complex data sample and is offloaded. HI uses a HI decision algorithm that acts on the on-device (local ML) model output for each input sample and decides if the sample is simple or complex. By only offloading the complex samples, HI reaps the benefits of performing inference on end devices without compromising on accuracy, and thus, it received considerable attention in the recent past [19, 20, 21, 15, 22, 23].
@@
Since HI requires inference on the device for every sample, it introduces a fixed overhead on latency and energy consumption per inference. Additionally, existing work on HI only studied the accuracy improvements over the on-device inference, without accounting for its latency and on-device energy consumption overheads. Furthermore, their conclusions were based only on simulations using abstract costs for offloading and classification errors. Finally, they ignore the diversity and heterogeneity of edge ML systems, including varying hardware, network connectivity, and models.
@@
In this paper, we address the aforementioned shortcomings and make multiple contributions. First, we perform accuracy, latency, and energy measurements for running embedded DL models on five devices with different capabilities and three image classification datasets (please refer to Section III for additional details about the devices and models used). These measurements will be of independent interest to the embedded ML community as these findings help optimize ML models for low-power, real-time applications, reducing reliance on cloud offloading while maintaining performance. In this respect, we have made the code for device measurements and model training publicly available. [^3] https://github.com/JoseGaDel/Hierarchical-Inference Second, we empirically evaluate the performance of HI and on-device inference and demonstrate that HI improves accuracy but has higher latency and energy consumption compared to on-device inference. This raises the question of the true benefits of using HI. Third, in contrast to the previous works, we argue that it is beneficial to choose an HI system that uses a smaller on-device model, even if the model does not meet the accuracy QoS requirement, rather than an HI system with a larger state-of-the-art on-device model that does. Through measurements, we demonstrate that an HI system with a smaller on-device model can achieve the accuracy QoS requirement while delivering lower latency and energy consumption. This highlights the advantages of a carefully designed HI system over standalone on-device inference. For instance, on the Raspberry Pi, with an accuracy requirement of $90$% for CIFAR-10 image classification, we demonstrate that using a smaller model, ResNet-8, with HI achieves $95.9$% accuracy while reducing latency by $73$% and energy consumption by $77$% compared to the larger on-device model, ResNet-56, which achieves $93.7$% accuracy. Additionally, HI delivers up to $11$% (absolute difference) higher accuracy for a given latency and energy requirement for the systems we studied. Fourth, for ImageNet-1K image classification, we found that offloading to remote inference remains the only viable solution for Coral Micro and Raspberry Pi, as no on-device models can meet the required low latency and energy consumption. To alleviate this issue, we propose using the early exit technique [24] to reduce the latency of inference on the device and consequently further improve the HI approach. In particular, we design a hybrid approach that combines Early Exit with HI (EE-HI) and demonstrate that, compared to HI, EE-HI reduces latency by up to $59.7$% and device energy consumption by up to $60.4$%.
@@
The rest of the paper is organized as follows. In Section II, we discuss the related works and the existing edge ML inference strategies. In Section III, we present the system model along with devices, datasets, models and performance metrics used in this study. In Section IV, we describe the system methodology used for measurements. We present the measurement results in Section V, and compare the performance of HI, on-device, and remote inference in Section VI. We present the design and the results of EE-HI in Section VII and conclude in Section VIII.
II Related Works
II-A Existing Strategies for Edge ML Inference
@@
This subsection discusses the related works in edge ML inference systems. The methods that use on-device inference can be broadly classified into 1) on-device inference, 2) DNN partitioning, 3) inference load balancing, and 4) HI. In the following, an inference task refers to generating the inference output for a given data sample at the device. For the on-device inference strategy, this task involves running the local DL model on the data sample. For remote ML inference, it entails offloading the data sample to a server and receiving the results. The HI task combines these approaches, either performing standalone on-device inference or using a combination of on-device and remote inference based on the HI offloading decision.
@@
On-Device Inference: As discussed in Section I, considerable effort has been dedicated to the design of compact DL models tailored for deployment on end devices [4, 25, 26, 27]. Further, techniques such as early exit [24] have been explored to reduce the execution time of DL inference. As a result, state-of-the-art compact DL models can perform simple ML inference tasks on devices, ranging from face recognition on mobile phones [28] to visual wake word and keyword spotting on IoT devices [8]. However, the main focus in designing these models has been improving accuracy, while time per inference and energy consumption have not been critically studied.
DNN Partitioning: DNN partitioning, proposed in [13], partitions the layers of a large-size DL model and deploys the front layers on the mobile device while the deep layers are deployed on an edge server. It received considerable attention recently [29, 30]. Recent research has even enhanced the performance of DNN partitioning by introducing techniques such as encoding and decoding data transmitted over networks [31] and selectively determining which features to process locally versus remotely [32]. But as demonstrated in [20], for DNN partitioning to be beneficial in reducing the execution time per inference, the processing times of the DL layers on the mobile device should be small relative to the communication time of the data generated between layers. Thus, this technique requires mobile GPUs, making it infeasible for resource-constrained end devices such as IoT devices.
Inference Load Balancing: Since the initial proposal of edge computing [33], significant attention has been devoted to the computational offloading for generic compute-intensive applications [34]. Due to the growing importance of edge intelligence, recent studies have examined computation offloading for applications using ML inference [35, 18, 14, 36]. This line of research aims to load-balance the inference task by partitioning the data samples between the device and the server, considering parameters such as job execution times, communication times, energy consumption, and average test accuracy of ML models. Nonetheless, in [37], the authors demonstrated that load-balancing data samples may lead to adversarial scenarios where most data samples scheduled on the device receive incorrect inferences. HI circumvents this issue by utilizing the local DL output and offloading a data sample if its local inference is likely incorrect.
@@
Hierarchical Inference: The key element is the HI decision algorithm for differentiating simple and complex samples, and so driving the offloading decision. To this end, all existing works use the soft-max values (confidence values) output by the local DL model corresponding to each class. An image is classified into the class that has the maximum soft-max value. Nikoloska and Zlatanov [19] computed the HI offloading decision by computing a threshold for the maximum soft-max value based on the transmission energy constraint of the device. Al-Atat et al. [20] proposed a general definition for HI, presented multiple use cases, and compared HI with existing distributed DL inference approaches at the edge. Similar to [19], a threshold was computed based on the trade-off between local misclassifications and offloading costs. Behera et al. [21] showed that using a binary LR on the first two highest soft-max values can improve the HI offloading decision. The binary LR is trained to learn which images are correctly classified and which are incorrectly classified by the local DL. Letsiou et al. [23] proposed a batched HI approach that offloads samples in batches to minimize communication overhead and enhance system efficiency while maintaining performance comparable to existing HI approaches. In this paper, we use the binary LR as the HI decision algorithm proposed in [21]. Nonetheless, all the aforementioned works use simulation to evaluate their performance against on-device inference using abstract costs for offloading and misclassification errors.
II-B Existing Measurement-based Studies
A key requirement for our work is to measure performance in order to evaluate the efficiency of ML models on different devices. However, measuring ML performance on resource-constrained and heterogeneous devices is challenging, as different hardware platforms typically require different software packages to run the models. To address this challenge, the ML performance (MLPerf) benchmarking organization has recently proposed a benchmarking suite, called TinyMLPerf, for tinyML models to perform simple inference tasks [8]. Throughput and latency measurements were presented in [38] for ResNet-50 and MobileNet on Jetson Nano and Coral Micro. Chu et al. [39] have proposed a more detailed and precise profiling method for Android devices through their tool, nnPerf, explicitly designed for TFLite models. Still, none of the above studies offer energy measurements or compare edge machine learning strategies.
An empirical comparison of on-device and cloud-based inference is presented in [40], illustrating the latency, energy, and accuracy trade-offs between the two approaches. The comparative study was limited to centralized techniques, leaving distributed inference mechanisms unexplored. Furthermore, some studies bench-marked diverse edge devices, such as the NVIDIA Jetson Nano (“end-of-life” as of March 2022) and Google Coral Dev Board, analyzing the on-device inference time [41], power consumption, and accuracy under varying workloads and model configurations [42]. Similarly to [40], these studies emphasized hardware-software performance comparisons without exploring the impact of distributed inference strategies. The authors in [43], conducted a measurement-based study exploring architectural trade-offs in deploying DNNs on edge devices with AI accelerators, focusing on model compression and DNN partitioning highlighting how these techniques balance energy efficiency, computational demands, and system performance, enabling edge devices with AI accelerators to support AI workloads effectively. However, as highlighted in [44], DNN partitioning is ineffective on devices lacking hardware accelerators, such as Raspberry Pi or MCUs. In contrast, this work presents a comprehensive measurement-based study evaluating the energy, latency, and accuracy trade-offs across on-device, remote, and HI approaches, covering a diverse range of edge devices from smaller MCUs to more capable platforms like the Jetson Orin. Furthermore, we also propose the EE-HI approach that reduces the latency and energy consumption of an HI system.
II-C Early Exit with various ML Inference
Early exit techniques [45, 46, 47] in DNNs optimize inference by enabling predictions at intermediate layers, thereby reducing computation and latency. BranchyNet [24] pioneered this approach, introducing confidence-based thresholds for early classification, specifically tailored for resource-constrained environments. MSDNet [48] advanced the concept by combining multi-scale feature extraction and early exits, allowing simpler inputs to exit early while processing more complex ones through deeper layers. EENet [49] introduced adaptive thresholds for dynamic, context-aware exits, while PTEENet [50] augmented pre-trained models with task-prioritized early exits for improved resource allocation. In [51], early exits are enhanced by self-distillation, guiding learning with intermediate layers to preserve accuracy. Innovations like EdgeExit [52] tailored early exits for heterogeneous edge systems, ensuring low-latency, power-efficient inference, while reinforcement learning-based methods [53] improved exit decisions through dynamic policy learning. These techniques are widely used in end devices for efficient, on-device inference.
Early exit has also been integrated with DNN partitioning methods to further enhance inference. In [54], DNNs are partitioned for efficient edge computing with early exits, optimizing resource usage on edge devices. Conversely, [55] proposes a hybrid method, partitioning DNNs across edge devices and the cloud, dynamically offloading computation as needed. Similar approaches in [17, 56, 57] demonstrate the effectiveness of combining early exits with DNN partitioning to improve inference efficiency. In contrast to previous works, we design a novel EE-HI system that uses the early exit technique to further reduce the latency and energy consumption of the HI system. More details of this method are presented in Section VII.
III ML Models and Performance Metrics
III-A Devices